# Radial and vertical constraints on the icy origin of HCO in the HD 163296 Protoplanetary Disk

### v1

@@ 4.2.1 Radial Distribution of HCO
From Figure 6, we deduce a clear correlation between the substructures of the profile with the CO snowline and , which is likely connected with changes in the chemistry. The same spatial coincidence observed in previous work of HCO toward HD 163296 (Carney et al., 2017; Guzmán et al., 2021) and other disks (e.g., Öberg et al., 2017; Kastner et al., 2018; Pegues et al., 2020) has been interpreted as an indicator of active HCO grain-surface production. Nevertheless, although less efficient at larger radii, the contribution from the gas-phase formation route may not be negligible.
@@
Despite warm temperatures ( K, see middle panel in Figure 6), there is a substantial amount of HCO in the inner disk ( au). Particularly, within the CO snowline, we would not expect to have the formation of abundant CO ices that could form HCO on dust grains. This warm HCO component is generally ascribed to warm gas-phase formation (Loomis et al., 2015; Öberg et al., 2017), but this could be complemented by an inward drift of ices with HCO already formed at slightly larger radii (Pegues et al., 2020).
@@
In the outer disk ( au), CO is expected to be abundant on ices, and therefore, the reaction represented by Equation 8 could become more efficient. In addition, the shielding from UV radiation is expected to decrease toward the outskirts of the disk, as suggested by previous DCO observations toward HD 163296 with a similar radial distribution (Salinas et al., 2017, 2018). Thus, the photodesorption of HCO (Carney et al., 2017; Öberg et al., 2017) and some CO (Öberg et al., 2015; Huang et al., 2016) from grains could increase, which could explain the bump of observed close to . However, the above can also be associated with an enhancement in the production of radicals and atomic O due to the photodissociation of CO in the upper layers, which would favor the gas-phase formation of HCO in the outer disk (Carney et al., 2017).
@@
In contrast to the and OPR profiles, does not present significant radial substructures, which could imply the dominance of a single formation mechanism throughout the disk, similar to was found in TW Hya (Terwisscha van Scheltinga et al., 2021). However, the temperature gradient in the inner disk is steeper than in the outer disk, rapidly decreasing from K to values around the disk-averaged K, and then decreases gently at larger radii. This distinction could support the previously suggested two-component HCO model, with a warm inner region with predominant gas-phase formation and a cold component beyond the CO snowline with contribution from both formation pathways (Loomis et al., 2015; Carney et al., 2017; Öberg et al., 2017; Pegues et al., 2020).
@@
Qi et al. (2015) estimated a CO freeze-out midplane temperature of K for HD 163296, which is a bit smaller than our derived value of K at au (the location of the CO snowline). Instead, typical CO freeze-out temperatures ( K, Öberg et al., 2015; Qi et al., 2015; Pinte et al., 2018) are reached in the outer disk, close to , where we expect to have more HCO from dust grains. The above is a consequence that disk temperature depends on the radius and the vertical height. Thus, even if the HCO emission comes from a radius outside the CO snowline, this same emission may have originated in a region where CO is not frozen out, emphasizing the importance of complementing our radial analysis with constraints on the vertical height.
4.2.2 Vertical Distribution of HCO
@@
We complemented our radial temperature measurements with the vertical distribution of the emission obtained with ALFAHOR, as shown in Figure 7. For consistency, we compared our vertical and radial constraints with predictions of the temperature structure in HD 163296 (see background maps in Figure 7). For the inner disk, values are more consistent with the empirical temperature structure (see left panel in Figure 7), while in the outer disk, resembles that predicted from thermo-chemical models (see right panel in Figure 7). These disparities align with the fact that thermo-chemical models can underpredict the temperature deduced from observations in the inner disk because they do not include mechanisms like accretion heating or CO depletion (Zhang et al., 2021). Moreover, in the outer disk, temperature predictions from CO observations could be more uncertain as the emission from optically thin CO isotopologues is less extended (Law et al., 2021b).
@@
Empirical measurements of the vertical height and predictions of the temperature structure agree that HCO is mostly tracing a cold region relatively close to the midplane (). The intersection of the CO snow surface ( K) with the surface is expected at au, depending on the assumed temperature structure (see dashdotted lines in Figure 7). This spatial feature is consistent with our delineation of the inner and outer disk, based on the column density profile, and could explain why the gas temperature traced by HCO has different behavior in these two regions, supporting different HCO formation pathways.
@@
Interestingly, the four analyzed transitions present modulations on the computed vertical height at similar positions as the integrated line intensity. Paneque-Carreño et al. (2023) resolved a similar structure in HD 163296 using the line but with higher angular resolution. They discarded the possibility of an effect of the line emission morphology in determining the vertical structure, linking it to actual changes in the vertical location of the HCO emitting region. The association of the local minimum heights with HCO emission rings (at and au, see Figure 6) could be related to an increased penetration of UV photons in the outer disk that would enhance the formation of HCO. Actually, vertical heights predicted in this work are near those layers with active photochemistry traced by HCN and CH (e.g., Law et al., 2021b; Paneque-Carreño et al., 2023). Still, the above would be consistent with the grain-surface chemistry (photodesorption of HCO), but also the cold gas-phase chemistry due to an increased production of radicals.
4.2.3 Ortho-to-para Ratio of HCO
@@
The OPR has been proposed as an indicator of the conditions under which HCO is formed, based on its relation with the spin temperature (, e.g., Guzmán et al., 2018). Owing to the large timescales involved in the ortho-para nuclear-spin conversion by non-reactive collisions in the gas-phase, is expected to resemble the formation temperature of HCO (Tudorie et al., 2006). According to the Boltzmann distribution, higher OPR values are associated with higher spin temperatures (see Guzmán et al., 2018, and references therein). Hence, a statistical value of was initially expected for the gas-phase chemistry, while lower values were thought to result when HCO forms on the surface of dust grains (Kahane et al., 1984; Guzmán et al., 2011). Nevertheless, the latter began to be questioned by laboratory experiments where water desorption from ice resulted in an OPR consistent with (Hama et al., 2016, 2018). Recent results suggest that the same behavior could occur for HCO desorption (Yocum et al., 2023), and previous work has discussed the possibility of low OPR values being associated with a cold gas-phase formation from radicals produced at low temperatures (Guzmán et al., 2018; Terwisscha van Scheltinga et al., 2021). It is, therefore, unclear how OPRs are related to the formation pathways of molecules.
@@
From our disk-averaged and resolved analysis, we derived a relatively constant value of OPR across the disk, which could still be consistent with 3.0, considering the uncertainties. Similar to the integrated intensity and the vertical height, the OPR seems to have some radial modulations that could be related to changes in the chemistry777We cannot discard the possibility that the tentative variations observed in the OPR profile could be due to slight deviations from the assumptions made to derive (e.g., LTE emission, same for both isomers and the same emitting region for all the lines).. The lowest OPR values are found close to the HCO rings and at lower disk altitudes; thus, they are likely associated with a rise in UV penetration as well. We also extracted the radial profile from the resolved OPR values (see Fig. 10), and find a coincidence between and values, mainly in the outer disk, which could indicate that the formation temperature of HCO is preserved. However, laboratory experiments have shown that does not correspond to the ice formation temperature when HCO is formed on ices by the UV photolysis of CHOH (Yocum et al., 2023). Still, it is also possible that could reflect the formation temperature if HCO forms more quiescently through the hydrogenation of CO ices.
@@
Therefore, our results are consistent with both the cold gas-phase and grain-surface formation scenarios in HD 163296. Indeed, more in-depth studies of dust-grain-related processes are needed to discern if grain-surface chemistry is critical for HCO formation in HD 163296, as well as its role in setting the OPR value. Since gas-phase products may also contribute to the formation of HCO on grains (Terwisscha van Scheltinga et al., 2021), chemical models considering both mechanisms is an alternative but, unfortunately, the physical parameters involved in grain-surface chemistry are poorly constrained (Guzmán et al., 2013). Modeling only gas-phase chemistry (e.g., Loomis et al., 2015, 2018a) could be a more straightforward approach, but it is beyond the scope of this work.
4.3 HCO in T Tauri versus Herbig Ae/Be disks
@@
Thanks to ALMA, HCO has been currently resolved in 20 protoplanetary disks (see Pegues et al., 2020; Terwisscha van Scheltinga et al., 2021, and references therein). From this sample, only five correspond to Herbig Ae/Be disks: the cold disks HD 163296 (this work) and MWC 480 (Pegues et al., 2020; Guzmán et al., 2021), and the warm disks IRS 48 (van der Marel et al., 2021; Booth et al., 2024a), HD 100546 (Booth et al., 2021, 2024b), and HD 169142 (Booth et al., 2023).
@@
MWC 480 has similar and values to HD 163296, while higher values and radial substructures of correlated with the dust continuum characterize IRS 48, HD 100546, and HD 169142. These differences reflect how HCO chemistry acts under different physical conditions since, typically, warm Herbig Ae/Be disks are characterized by dust cavities, which prevent the in-situ formation of CO-ices, and by the presence of CHOH detections (Booth et al., 2023, 2024a, 2024b). In these cases, HCO formation is probably dominated by warm gas-phase chemistry and/or thermal desorption from inherited ices, as suggested by the low OPR of deduced in IRS 48 (van der Marel et al., 2021).
@@
Instead, the morphology of and the disk-averaged of HCO in cold Herbig Ae/Be disks are more similar to those observed in T Tauri disks (e.g., Pegues et al., 2020). Still, the values of cold Herbig Ae/Be disks are systematically lower than those of T Tauri disks, which has been explained as a consequence of the higher temperatures expected around more massive stars that limits the CO freeze-out (Pegues et al., 2020), in line with the two-component HCO model discussed in Section 4.2.2. To verify this possibility of a similar chemistry, we can analyze the radial behavior of the OPR, where the only comparison that can be made is with the T Tauri disk TW Hya. Interestingly, unlike the relatively constant OPR observed in HD 163296, the OPR profile of TW Hya has an apparent decreasing behavior (Terwisscha van Scheltinga et al., 2021), thus suggesting a different formation process. Further experiments of the OPR in HCO ices and more determinations of the OPR around both T Tauri and Herbig Ae/Be protoplanetary disks are needed to confirm this scenario.
@@
TW Hya is a relatively peculiar source. It is relatively old compared to other T Tauri stars, and it is the only T Tauri disk with known detections of CHOH (Walsh et al., 2016). Also, in terms of HCO, it has a relatively high and nearly constant excitation temperature ( K) and a different column density morphology (no rings in the outer disk), in comparison with other T Tauri disks and HD 163296. Moreover, constraints on the vertical structure have placed the emitting surface of HCO in TW Hya at higher heights (, Öberg et al., 2017; Terwisscha van Scheltinga et al., 2021) than in other disks with empirical determinations (, Paneque-Carreño et al., 2023). These differences align with the fact that TW Hya is one of the few disks where HCO is proposed to be predominantly formed by gas-phase rather than a combination of both formation pathways (Terwisscha van Scheltinga et al., 2021). Observations toward other disks are needed to better understand the dominant formation pathway of HCO.
@@
Curiously, even though it has no efficient gas-phase formation pathways, CHOH is relatively abundant in TW Hya (CHOH/HCO ) but not so in HD 163296 (CHOH/HCO ), according to the upper limits derived by Carney et al. (2019). One possibility to address this counterintuitive behavior could be a difference in the UV radiation field. UV photons can photodesorb intact CHOH from ices, but also photofragments of HCO (Guzmán et al., 2013; Bertin et al., 2016; Yocum et al., 2023). Because massive Herbig Ae/Be disks have stronger UV fluxes than T Tauri disks (Walsh et al., 2015), maybe CHOH tends to be more fragmented for sources like HD 163296, thus contributing to HCO emission instead of the gas-phase CHOH, as supposed in TW Hya.
@@
Furthermore, laboratory experiments have demonstrated that X-ray photodesorption is also an efficient mechanism to release CHOH into the gas-phase, even more than UV photodesorption (Basalgète et al., 2021a, b). The above could also contribute to differences in the CHOH/HCO ratio since T Tauri sources have larger X-ray luminosities than Herbig Ae/Be (Ryspaeva et al., 2023) and, therefore, the X-ray-induced chemistry is predicted to be more significant (Walsh et al., 2015). More complementary detections of HCO and upper limits for CHOH (if not detected) in protoplanetary disks are needed to clarify this behavior.

### v2

@@ 4.2.1 Radial Distribution of H2CO
From Figure 6, we deduce a clear correlation between the substructures of the profile with the CO snowline and , which is likely connected with changes in the chemistry. The same spatial coincidence observed in previous work of H2CO toward HD 163296 (Carney et al., 2017; Guzmán et al., 2021) and other disks (e.g., Öberg et al., 2017; Kastner et al., 2018; Pegues et al., 2020) has been interpreted as an indicator of active H2CO grain-surface production. Nevertheless, although less efficient at larger radii, the contribution from the gas-phase formation route may not be negligible.
@@
Despite warm temperatures ( K, see middle panel in Figure 6), there is a substantial amount of H2CO in the inner disk ( au). Particularly, within the CO snowline, we would not expect to have the formation of abundant CO ices that could form H2CO on dust grains. This warm H2CO component is generally ascribed to warm gas-phase formation (Loomis et al., 2015; Öberg et al., 2017), but this could be complemented by an inward drift of ices with H2CO already formed at slightly larger radii (Pegues et al., 2020).
@@
In the outer disk ( au), CO is expected to be abundant on ices, and therefore, the reaction represented by Equation 8 could become more efficient. In addition, the shielding from UV radiation is expected to decrease toward the outskirts of the disk, as suggested by previous DCO+ observations toward HD 163296 with a similar radial distribution (Salinas et al., 2017, 2018). Thus, the photodesorption of H2CO (Carney et al., 2017; Öberg et al., 2017) and some CO (Öberg et al., 2015; Huang et al., 2016) from grains could increase, which could explain the bump of observed close to . However, the above can also be associated with an enhancement in the production of radicals and atomic O due to the photodissociation of CO in the upper layers, which would favor the gas-phase formation of H2CO in the outer disk (Carney et al., 2017).
@@
In contrast to the and OPR profiles, does not present significant radial substructures, which could imply the dominance of a single formation mechanism throughout the disk, similar to was found in TW Hya (Terwisscha van Scheltinga et al., 2021). However, the temperature gradient in the inner disk is steeper than in the outer disk, rapidly decreasing from K to values around the disk-averaged K, and then decreasing gently at larger radii. This distinction could support the previously suggested two-component H2CO model, with a warm inner region with predominant gas-phase formation and a cold component beyond the CO snowline with contribution from both formation pathways (Loomis et al., 2015; Carney et al., 2017; Öberg et al., 2017; Pegues et al., 2020).
@@
Qi et al. (2015) estimated a CO freeze-out midplane temperature of K for HD 163296, which is a bit smaller than our derived value of K at au (the location of the CO snowline). Instead, typical CO freeze-out temperatures ( K, Öberg et al., 2015; Qi et al., 2015; Pinte et al., 2018) are reached in the outer disk, close to , where we expect to have more H2CO from dust grains. The above is a consequence that disk temperature depends on the radius and the vertical height. Thus, even if the H2CO emission comes from a radius outside the CO snowline, this same emission may have originated in a region where CO is not frozen out, emphasizing the importance of complementing our radial analysis with constraints on the vertical height.
4.2.2 Vertical Distribution of H2CO
@@
We complemented our radial temperature measurements with the vertical distribution of the emission obtained with ALFAHOR, as shown in Figure 7. For consistency, we compared our vertical and radial constraints with predictions of the temperature structure in HD 163296 (see background maps in Figure 7). For the inner disk, values are more consistent with the empirical temperature structure (see left panel in Figure 7), while in the outer disk, resembles that predicted from thermochemical models (see right panel in Figure 7). These disparities align with the fact that thermochemical models can underpredict the temperature deduced from observations in the inner disk because they do not include mechanisms like accretion heating or CO depletion (Zhang et al., 2021). Moreover, in the outer disk, temperature predictions from CO observations could be more uncertain as the emission from optically thin CO isotopologues is less extended (Law et al., 2021b).
@@
Empirical measurements of the vertical height and predictions of the temperature structure agree that H2CO is mostly tracing a cold region relatively close to the midplane (). The intersection of the CO snow surface ( K) with the surface is expected at au, depending on the assumed temperature structure (see dash-dotted lines in Figure 7). This spatial feature is consistent with our delineation of the inner and outer disk, based on the column density profile, and could explain why the gas temperature traced by H2CO has different behavior in these two regions, supporting different H2CO formation pathways.
@@
Interestingly, the four analyzed transitions present modulations on the computed vertical height at similar positions as the integrated line intensity. Paneque-Carreño et al. (2023) resolved a similar structure in HD 163296 using the line but with higher angular resolution. They discarded the possibility of an effect of the line emission morphology in determining the vertical structure, linking it to actual changes in the vertical location of the H2CO emitting region. The association of the local minimum heights with H2CO emission rings (at and au, see Figure 6) could be related to an increased penetration of UV photons in the outer disk that would enhance the formation of H2CO. Actually, vertical heights predicted in this work are near those layers with active photochemistry traced by HCN and C2H (e.g., Law et al., 2021b; Paneque-Carreño et al., 2023). Still, the above would be consistent with the grain-surface chemistry (photodesorption of H2CO), but also the cold gas-phase chemistry due to increased production of radicals.
4.2.3 Ortho-to-para Ratio of H2CO
@@
The OPR has been proposed as an indicator of the conditions under which H2CO is formed, based on its relation with the spin temperature (, e.g., Guzmán et al., 2018). Owing to the large timescales involved in the ortho-para nuclear-spin conversion by nonreactive collisions in the gas phase, is expected to resemble the formation temperature of H2CO (Tudorie et al., 2006). According to the Boltzmann distribution, higher OPR values are associated with higher spin temperatures (see Guzmán et al., 2018, and references therein). Hence, a statistical value of was initially expected for the gas-phase chemistry, while lower values were thought to result when H2CO forms on the surface of dust grains (Kahane et al., 1984; Guzmán et al., 2011). Nevertheless, the latter began to be questioned by laboratory experiments where water desorption from ice resulted in an OPR consistent with (Hama et al., 2016, 2018). Recent results suggest that the same behavior could occur for H2CO desorption (Yocum et al., 2023), and previous work has discussed the possibility of low OPR values being associated with a cold gas-phase formation from radicals produced at low temperatures (Guzmán et al., 2018; Terwisscha van Scheltinga et al., 2021). It is, therefore, unclear how OPRs are related to the formation pathways of molecules.
@@
From our disk-averaged and resolved analysis, we derived a relatively constant value of OPR across the disk, which could still be consistent with 3.0, considering the uncertainties. Similar to the integrated intensity and the vertical height, the OPR seems to have some radial modulations that could be related to changes in the chemistry777We cannot discard the possibility that the tentative variations observed in the OPR profile could be due to slight deviations from the assumptions made to derive (e.g., LTE emission, same for both isomers and the same emitting region for all the lines).. The lowest OPR values are found close to the H2CO rings and at lower disk altitudes; thus, they are likely associated with a rise in UV penetration as well. We also extracted the radial profile from the resolved OPR values (see Fig. 10), and found a coincidence between and values, mainly in the outer disk, which could indicate that the formation temperature of H2CO is preserved. However, laboratory experiments have shown that does not correspond to the ice formation temperature when H2CO is formed on ices by the UV photolysis of CH3OH (Yocum et al., 2023). Still, it is also possible that could reflect the formation temperature if H2CO forms more quiescently through the hydrogenation of CO ices.
@@
Therefore, our results are consistent with both the cold gas-phase and grain-surface formation scenarios in HD 163296. Indeed, more in-depth studies of dust-grain-related processes are needed to discern if grain-surface chemistry is critical for H2CO formation in HD 163296, as well as its role in setting the OPR value. Since gas-phase products may also contribute to the formation of H2CO on grains (Terwisscha van Scheltinga et al., 2021), chemical models considering both mechanisms are an alternative but, unfortunately, the physical parameters involved in grain-surface chemistry are poorly constrained (Guzmán et al., 2013). Modeling only gas-phase chemistry (e.g., Loomis et al., 2015, 2018a) could be a more straightforward approach, but it is beyond the scope of this work.
4.3 H2CO in T Tauri versus Herbig Ae/Be disks
@@
Thanks to ALMA, H2CO has been currently resolved in 20 protoplanetary disks (see Pegues et al., 2020; Terwisscha van Scheltinga et al., 2021, and references therein). From this sample, only five correspond to Herbig Ae/Be disks: the cold disks HD 163296 (this work) and MWC 480 (Pegues et al., 2020; Guzmán et al., 2021), and the warm disks IRS 48 (van der Marel et al., 2021; Booth et al., 2024a), HD 100546 (Booth et al., 2021, 2024b), and HD 169142 (Booth et al., 2023).
@@
MWC 480 has similar and values to HD 163296, while higher values and radial substructures of correlated with the dust continuum characterize IRS 48, HD 100546, and HD 169142. These differences reflect how H2CO chemistry acts under different physical conditions since, typically, warm Herbig Ae/Be disks are characterized by dust cavities, which prevent the in situ formation of CO ices, and by the presence of CH3OH detections (Booth et al., 2023, 2024a, 2024b). In these cases, H2CO formation is probably dominated by warm gas-phase chemistry and/or thermal desorption from inherited ices, as suggested by the low OPR of deduced in IRS 48 (van der Marel et al., 2021).
@@
Instead, the morphology of and the disk-averaged of H2CO in cold Herbig Ae/Be disks are more similar to those observed in T Tauri disks (e.g., Pegues et al., 2020). Still, the values of cold Herbig Ae/Be disks are systematically lower than those of T Tauri disks, which has been explained as a consequence of the higher temperatures expected around more massive stars that limit the CO freeze-out (Pegues et al., 2020), in line with the two-component H2CO model discussed in Section 4.2.2. To verify this possibility of similar chemistry, we can analyze the radial behavior of the OPR, where the only comparison that can be made is with the T Tauri disk TW Hya. Interestingly, unlike the relatively constant OPR observed in HD 163296, the OPR profile of TW Hya has an apparent decreasing behavior (Terwisscha van Scheltinga et al., 2021), thus suggesting a different formation process. Further experiments of the OPR in H2CO ices and more determinations of the OPR around both T Tauri and Herbig Ae/Be protoplanetary disks are needed to confirm this scenario.
@@
TW Hya is a relatively peculiar source. It is relatively old compared to other T Tauri stars, and it is the only T Tauri disk with known detections of CH3OH (Walsh et al., 2016). Also, in terms of H2CO, it has a relatively high and nearly constant excitation temperature ( K) and a different column density morphology (no rings in the outer disk), in comparison with other T Tauri disks and HD 163296. Moreover, constraints on the vertical structure have placed the emitting surface of H2CO in TW Hya at higher heights (, Öberg et al., 2017; Terwisscha van Scheltinga et al., 2021) than in other disks with empirical determinations (, Paneque-Carreño et al., 2023). These differences align with the fact that TW Hya is one of the few disks where H2CO is proposed to be predominantly formed by gas phase rather than a combination of both formation pathways (Terwisscha van Scheltinga et al., 2021). Observations toward other disks are needed to better understand the dominant formation pathway of H2CO.
@@
Curiously, even though it has no efficient gas-phase formation pathways, CH3OH is relatively abundant in TW Hya (CH3OH/H2CO ) but not so in HD 163296 (CH3OH/H2CO ), according to the upper limits derived by Carney et al. (2019). One possibility to address this counterintuitive behavior could be a difference in the UV radiation field. UV photons can photodesorb intact CH3OH from ices, but also photofragments of H2CO (Guzmán et al., 2013; Bertin et al., 2016; Yocum et al., 2023). Because massive Herbig Ae/Be disks have stronger UV fluxes than T Tauri disks (Walsh et al., 2015), maybe CH3OH tends to be more fragmented for sources like HD 163296, thus contributing to H2CO emission instead of the gas-phase CH3OH, as supposed in TW Hya.
@@
Furthermore, laboratory experiments have demonstrated that X-ray photodesorption is also an efficient mechanism to release CH3OH into the gas phase, even more than UV photodesorption (Basalgète et al., 2021a, b). The above could also contribute to differences in the CH3OH/H2CO ratio since T Tauri sources have larger X-ray luminosities than Herbig Ae/Be (Ryspaeva et al., 2023), and therefore, the X-ray-induced chemistry is predicted to be more significant (Walsh et al., 2015). More complementary detections of H2CO and upper limits for CH3OH (if not detected) in protoplanetary disks are needed to clarify this behavior.